OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and third equations, then the second and third equations.

OpenStudy (anonymous):

x – y + 2z = –2 2x + 2y + z = 7 3x + 3y – z = 3

OpenStudy (anonymous):

choices are:

OpenStudy (anonymous):

2x + 2y = 7 3x + 3y = 3 2x + 2y = 7 5x + 5y = 10 7x + 5y = 4 3x + 3y = 3 7x + 5y = 4 5x + 5y = 10

OpenStudy (anonymous):

for first and third equations...: by what do you multiply the third equation in order to eliminate z when the two equations are added?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

yes, and what's the result when you add them after multiplying the third eqn. by 2?

OpenStudy (anonymous):

-5x - 7y = -8

OpenStudy (anonymous):

I got C, is that right?

OpenStudy (anonymous):

x – y + 2z = –2 +2(3x + 3y – z = 3) ------------------

OpenStudy (anonymous):

x - y + 2z = -2 6x + 6y -2z = 6

OpenStudy (anonymous):

so 7x + 5y = 4

OpenStudy (anonymous):

so it's C or D now add equations 2 and 3

OpenStudy (anonymous):

so it's either c or d hmm

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

5x + 5y = 10

OpenStudy (anonymous):

it's d?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

sure